# Reliably Detecting Model Failures in Deployment Without Labels

### v3

@@ C.5 Comparisons with a fully Bayesian network
Figure 9: Toy experiment on the Two Moons dataset generator Pedregosa et al. [2011]. $500$ points are noisily sampled from each class at noise level $0.3$. We visualize $200$ total samples here for clarity. (Figure 9) A fully Bayesian neural network is trained on this dataset. The decision boundaries correspond to the level set of the probabilities of predicting either class at level $0.5$. Forward passing the full grid through this net, we find different decision boundaries pictured in green. (Figure 9) A comparable VBLL model is trained. Without our double sampling scheme, most decision boundaries samples collapse onto the mean model’s boundary, highlighted by the yellow region. Figures 9, 9, and 9 depict probabilities under double sampling with temperatures $\tau=2,5,10$ respectively. We observe that sampling predictions from gradually softer probabilities give noisier estimates across the space. Although this does not provide a concrete continuous level set at $0.5$, the colors depict the distribution of predictions from which D3M samples, which “simulates” the diversity of decision boundaries.

### v4

@@ C.5 Comparisons with a fully Bayesian network
Figure 9: Toy experiment on the Two Moons dataset generator [69]. $500$ points are noisily sampled from each class at noise level $0.3$. We visualize $200$ total samples here for clarity. (Figure 9) A fully Bayesian neural network is trained on this dataset. The decision boundaries correspond to the level set of the probabilities of predicting either class at level $0.5$. Forward passing the full grid through this net, we find different decision boundaries pictured in green. (Figure 9) A comparable VBLL model is trained. Without our double sampling scheme, most decision boundaries samples collapse onto the mean model’s boundary, highlighted by the yellow region. Figures 9, 9, and 9 depict probabilities under double sampling with temperatures $\tau=2,5,10$ respectively. We observe that sampling predictions from gradually softer probabilities give noisier estimates across the space. Although this does not provide a concrete continuous level set at $0.5$, the colors depict the distribution of predictions from which D3M samples, which “simulates” the diversity of decision boundaries.